(1) diamond is one form of elemental carbon. An engagement right contains diamond weigh 1.25 carat. How many atoms are present in the diamond?


(2) A 55 kg woman has 7.5×10^-3 mol of haemoglobin in her blood. How many haemoglobin in this? What is this quantity in gram?


(3) A 24.81 sample of gaseous compound containing only carbon, oxygen and chlorine is determined to contain 3.01gram of carbon, 4.0 gram of oxygen and 17.81 gram of chlorine. What is the compound's percent composition?


(4) Determine the empirical and molecular formula for chrysotle asbestos, chrysotle has the following percent composition.

28.03% magnesium, 21.60% silicon, 1.16% hydrogen and 49.21% of oxygen.


1
Expert's answer
2021-03-02T04:56:12-0500

(1) 1.25 carat = 0.2*1.25 = 0.25 g. n(C) = m/M = (0.25 g)/(12 g/mol) = 0.02083 mol

N(C) = n(C)*NA = (0.02083 mol)*(6.02*1023 mol-1) = 1.254*1022


(2) 7.5×10-3 mol of haemoglobin is (7.5×10-3 mol)*(6.02*1023 mol-1) = 4.515*1021 molecules

Haemoglobin molar mass is approximately 65 000 g/mol

m(haemoglobin) = (7.5×10-3 mol)*(65 000 g/mol) = 487.5 g


(3) w(C) = 3.01/24.81*100% = 12.1%

w(O) = 4/24.81*100% = 16.1%

w(Cl) = 17.81/24.81*100% = 71.8%


(4) 28.03% magnesium, 21.60% silicon, 1.16% hydrogen and 49.21% of oxygen by mass

28.03/24 = 1.168 magnesium, 21.60/28 = 0.771 silicon, 1.16 hydrogen and 49.21/16 = 3.076 of oxygen by moles

Mg:Si:H:O = 1.168:0.771:1.16:3.076 = 1.5:1:1.5:4 = 3:2:3:8

empirical formula Mg3Si2H3O8

72.915 + 56.172 + 3.024 + 127.992 = 260.103 g/mol Mg3Si2H3O8

molar mass of chrysotle asbestos is 520.8 g/mol

520.8 g/mol. / 260.103 g/mol = 2

2*(Mg3Si2H3O8) = Mg6Si4H6O16

molecular formula Mg6Si4H6O16